### v2

@@ Abstract.
Conformal prediction (CP) is an Uncertainty Representation technique that delivers finite-sample calibrated prediction regions for any underlying Machine Learning model. Its status as an Uncertainty Quantification (UQ) tool, though, has remained conceptually opaque: While Conformal Prediction Regions (CPRs) give an ordinal representation of uncertainty (larger regions typically indicate higher uncertainty), they lack the capability to cardinally quantify it (twice as large regions do not imply twice the uncertainty). We adopt a category-theoretic approach to CP—framing it as a morphism, embedded in a commuting diagram, of two newly-defined categories—that brings us three joys. First, we show that—under minimal assumptions—CP is intrinsically a UQ mechanism, that is, its cardinal UQ capabilities are a structural feature of the method. Second, we demonstrate that CP bridges (and perhaps subsumes) the Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning. Finally, we show that a CPR is the image of a covariant functor. This observation is relevant to AI privacy: It implies that privacy noise added locally does not break the global coverage guarantee.
@@ 1. Introduction
For example, Stehlík, (2012) and Song, (2024) survey how to frame statistical learning models into a CT framework. Shiebler et al., (2021) do the same for gradient-based methods, probabilistic methods, and invariant and equivariant learning. Furthermore, Cornish, (2025) expands on the latter by studying the problem of symmetrizing a neural network along a group homomorphism. That is, given two groups $H$ and $G$, they study a procedure that converts $H$-equivariant neural networks to $G$-equivariant ones. More in general, CT approaches to probability and statistics (which play a pivotal role in modern-day ML) are active fields of study, see e.g. Fritz et al., (2025); Bohinen and Perrone, (2025); Ensarguet and Perrone, (2024); Ackerman et al., (2024); Matache et al., (2022), and are widely discussed in the nLab and in the n-Category Café.
@@
    • As a byproduct, the user will now be able to see (i) how the choice of non-conformity measure $\psi$ controls the size of the prediction region, and (ii) how much it affects (different types of) uncertainty in the analysis.
@@
  • Privacy by Design. Seeing the Conformal Prediction Region as the image of a covariant functor directly benefits AI privacy (Fang et al.,, 2025): (i) Any locally added privacy noise leaves coverage guarantees intact, and (ii) In federated learning (Liu et al.,, 2024), each agent need only share a summary object (the credal set $\mathcal{M}$, introduced in the following paragraphs) instead of raw data.
@@
To make the paper largely self-contained, we give the necessary background on CP in Section 2. There, we also provide some concepts from the Imprecise Probability (IP) literature. IPs (Walley,, 1991; Augustin et al.,, 2014; Troffaes and de Cooman,, 2014) are mathematical models that allow the scholar to reason in the presence of uncertainty (as for classical probabilities), and ambiguity (that is, ignorance, full or partial, about the “true” data generating process). [^4] For more modern references, see e.g. (Caprio and Gong,, 2023; Caprio and Mukherjee,, 2023; Caprio and Seidenfeld,, 2023; Caprio et al., 2024b, ; Caprio,, 2024, 2025). Their applications to ML form a lively subject field, see e.g. Caprio et al., 2024a ; Sale et al., (2023, 2024); Caprio et al., 2024c ; Lu et al., (2024); Dutta et al., (2024); Chau et al., (2025). Functional Analysis and Category Theory notions are brushed up in Appendices A and B, respectively. The structure of the paper can be summarized as follows.
@@
In Sections 3 and 4, we introduce two categories, $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$, respectively. For both, the objects are topological spaces; The morphisms are, respectively, upper hemicontinuous, and weakly measurable correspondences. We study these categories because continuity and measurability are fundamental—but generally non-equivalent—concepts in functional analysis. Under minimal assumptions on the state space $Y$, on the choice of the credibility level $\alpha$, and on the class $\mathscr{F}$ of non-conformity scores $\psi$, we show that CP is a morphism of both. Viewing CP as a morphism rather than a stand-alone algorithm lets ML engineers plug it into functional-style libraries (e.g. JAX or PyTorch Functorch) and obtain automatic batching, JIT compilation, and differentiability of the score map, while the coverage guarantee remains mathematically intact.
@@
Not only is the CP correspondence a morphism of these categories, but it is also part of a diagram that commutes. As we shall see (in particular in Example 1 and in the discussion thereafter), this discovery implies the intrinsic (cardinal) UQ capabilities of CP.
@@
In Section 5, we show that, under minimal assumption, a diagram that codifies the way Bayesian CP (Fong and Holmes,, 2021), classical CP, and CP-informed imprecise probability theory generate an $\alpha$-level prediction region, commutes. This is a strong indicator that CP forms a bridge between, and perhaps subsumes, the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning (Martin,, 2022).
@@
Section 6 expands on category-theoretic Imprecise Probability theory (Liell-Cock and Staton,, 2024). We formalize the notion of Imprecise Highest Density Region (Coolen,, 1992)—that can be thought of as a “robust version” of a confidence interval which also accounts for distribution uncertainty—as a (covariant) functor $\mathsf{IHDR}_{\alpha}$, $\alpha\in[0,1]$, between the category of closed and convex sets of finitely additive probabilities on $Y$, and that of subsets of $Y$, both with inclusion morphisms. In turn, the Conformal Prediction Region (CPR) can be seen as the image $\mathsf{IHDR}_{\alpha}(\mathcal{M})$ of such a functor at the set of probabilities $\mathcal{M}$ associated with the CP procedure (Cella and Martin,, 2022). This has e.g. the following promising implication. If we have many sources producing credal sets $\mathcal{M}_{j}$, any monotone, measurable or privacy-preserving operation applied locally to each credal set $\mathcal{M}_{j}$ propagates automatically through the functor. The resulting prediction regions, then, still satisfy the global coverage guarantee while only the (possibly privatized) summary objects $\mathcal{M}_{j}$, never the raw data, need be shared.
@@ 2.1. Full Conformal Prediction
This section is an expanded reproduction of Caprio et al., 2025a (, Section 2.1). It is based on Cella and Martin, (2022, Section 4.1), who summarize work by Shafer and Vovk, (2008) and Vovk et al., (2005) on Transductive (of Full) Conformal Prediction.
@@
for all $i\in\{1,\ldots,n+1\}$, where $y^{n+1}_{-i}=y^{n+1}\setminus\{y_{i}\}$ and $\psi:Y^{n}\times Y\rightarrow\mathbb{R}$ is a fixed function that is invariant to permutations in its first vector argument. Function $\psi$, which is called a non-conformity measure, is constructed in such as way that $\psi_{i}(y^{n+1})$ is small if and only if $y_{i}$ agrees with—i.e. is “close to”—a prediction based on the data $y^{n+1}_{-i}$. Or, the other way around, large values $\psi_{i}(y^{n+1})$ suggest that the observation $y_{i}$ is “strange” and does not conform to the rest of the data $y^{n+1}_{-i}$. The key idea is to define $\psi_{i}(y^{n+1})$ in a way that allows one to compare $y_{i}$ to a suitable summary of $y^{n+1}_{-i}$, e.g.  $\psi_{i}(y^{n+1})=|\text{mean}(y^{n+1}_{-i})-y_{i}|$, for all $i\in\{1,\ldots,n+1\}$. Notice that transformation $\mathcal{Y}^{n+1}\rightarrow T^{n+1}$ preserves exchangeability.
@@ 2.2. Imprecise Probabilities
Let $(Y,\Sigma_{Y})$ be a measurable space, and call $\Delta_{Y}$ the space of finitely additive probabilities on $Y$, endowed with the weak topology. A credal set $\mathcal{M}$ is a weak-closed and convex element of $\Delta_{Y}$ (Levi,, 1980). Throughout the paper, we will denote by $\mathscr{C}\subset 2^{\Delta_{Y}}$ the space of nonempty weak-closed and convex subsets of $\Delta_{Y}$. For any non-empty credal set $\mathcal{M}\in\mathscr{C}$, define the upper probability as the upper envelope of $\mathcal{M}$,
@@
Even in the presence of the latter, we are able to define (a version of) a confidence interval. Fix a credibility level $\alpha\in[0,1]$. Then, the $\alpha$-level Imprecise Highest Density Region (IHDR) of credal set $\mathcal{M}$ is the set $B_{\mathcal{M}}^{\alpha}\subset Y$ such that $\underline{P}(B_{\mathcal{M}}^{\alpha})=1-\alpha$, and its “size is minimal” (Coolen,, 1992). More formally (Caprio et al., 2025a, , Equation (11)),
@@
We also consider a slight modification of $\mathsf{CRED}$, that will significantly simplify our proving endeavors in Sections 3, 4, and 5; We denote it by $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\text{, }\forall A\in\Sigma_{Y}\text{ closed}\}$. As we can see, $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\supseteq\mathcal{M}(\overline{\Pi})$, because it imposes less constraints on the $P$’s: They only have to be setwise dominated on closed sets. The upper probability $\overline{\Pi}$ and its conjugate lower probability $\underline{\Pi}$ (the only characteristics of the credal sets that we use) stay the same for both $\mathsf{CRED}$ and $\mathsf{CRED}_{\text{cl}}$: This justifies our simplifying choice. To avoid notation complications, in the remainder of this work we do not distinguish between $\mathsf{CRED}(y^{n},\psi)$ and $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)$.
@@ 3. Conformal Prediction as an Upper Hemicontinuous Morphism
  • Morphisms: Upper hemicontinuous correspondences $\Phi:X\rightrightarrows Y$ (see Definition 26 in Appendix A).
@@
We verify the axioms of a category.
@@
(3) Closure under composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be upper hemicontinuous correspondences. By Aliprantis and Border, (2006, Theorem 17.23), the composition $\Psi\circ\Phi$ defined by
$(\Psi\circ\Phi)(x)\coloneqq\bigcup_{y\in\Phi(x)}\Psi(y)$
is also upper hemicontinuous. Thus, the composition of morphisms is again a morphism.
@@
    $\Delta_{Y}$ is endowed with the weak topology, and $\mathscr{C}\subset 2^{\Delta_{Y}}$ is endowed with the Vietoris topology (see Definition 29 in Appendix A).
@@
Another important lemma that sets the table for our main result is the following.
@@
Proof.
$\pi_{\psi}(y_{n+1},y^{n})\;\geq\;\frac{k}{n+1}.$
Step 4 (Take the Supremum). Because $y_{n+1}\in A$,
@@
Fix any $n\in\mathbb{N}$. We first show that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ have a closed graph.
(i) $\kappa$. By (ii) and its definition, the conformal transducer $\pi(\cdot,y^{n},\psi):Y\to[0,1]$ is upper semicontinuous. Because $\alpha$ avoids the grid $S_{n+1}$, let
@@
This is a super-level set of $\pi(\cdot,y^{n},\psi)$, and since $\pi$ is upper semicontinuous, $G$ is closed. In turn, $\operatorname{Gr}(\kappa)=G$ is closed.
@@
Weak-continuity of $P\mapsto P(A)$ and upper semicontinuity of the right-hand side (see Lemma 4) make $G_{A}$ closed. Because $\operatorname{Gr}(\mathsf{CRED})=\bigcap_{A\in\Sigma_{Y}}G_{A}$, the graph is closed.
(iii) $\mathsf{IHDR}$. For each $A\in\Sigma_{Y}$ define $f_{A}(\alpha,\mathcal{M})\coloneqq\underline{P}_{\!\mathcal{M}}(A)-(1-\alpha)$. Weak continuity of $P\mapsto P(A)$ and compactness of $\mathcal{M}$ imply that $\mathcal{M}\mapsto\underline{P}_{\!\mathcal{M}}(A)$ is lower semicontinuous in the Vietoris topology (iv), while $\alpha\mapsto 1-\alpha$ is continuous. Hence the set
$H_{A}\coloneqq\bigl{\{}(\alpha,\mathcal{M}):f_{A}(\alpha,\mathcal{M})\geq 0\bigr{\}}$
is closed; Moreover $(\alpha,\mathcal{M})\mapsto H_{A}(\alpha,\mathcal{M})$ is closed-valued by (i). In turn, $\operatorname{Gr}(\mathsf{IHDR})=\cap_{A\in\Sigma_{Y}}(([0,1]\times\mathscr{C})\setminus H_{A})\times Y\cup(H_{A}\times A)$ is closed.
Notice also that $Y$ is compact Hausdorff by (i), and that $\Delta_{Y}$ is compact by (iv) and Banach-Alaoglu, and Hausdorff because it is a subset of a dual Banach space equipped with its weak topology.
It follows that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are upper hemicontinuous by Lemma 3. In turn, they are morphisms of $\mathbf{UHCont}$. Fix now some $\alpha\in[0,1]\setminus S_{n+1}$. The restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ remain upper hemicontinuous. Notice also that $t(\mathsf{CRED})=s(\mathsf{IHDR}_{\alpha})=\mathscr{C}$; That is, the target space of $\mathsf{CRED}$ and the source space of $\mathsf{IHDR}_{\alpha}$ coincide, and are equal to the space of credal sets $\mathscr{C}$. This is clear by looking at the functional forms $\mathsf{CRED}^{\text{funct}}$ and $\mathsf{IHDR}^{\text{funct}}_{\alpha}$ of $\mathsf{CRED}$ and $\mathsf{IHDR}_{\alpha}$, respectively.
$Y$ is a topological space from (i); $[0,1]\times\mathscr{C}$ is a topological space in the product topology by (iv); $Y^{n}\times\mathscr{F}$ is a topological space in the product topology by (ii); $[0,1]\times Y^{n}\times\mathscr{F}$ is a topological space in the product topology by (i) and (ii). Hence, these spaces are all objects of $\mathbf{UHCont}$.
The fact that the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}$, then, is an immediate consequence of Caprio et al., 2025a (, Proposition 5). It states that, for all $\alpha\in[0,1]$, $y^{n}\in Y^{n}$, and $\psi\in\mathscr{F}$, the Conformal Prediction Region $\kappa_{\alpha}(y^{n},\psi)$ corresponds to the $\alpha$-level IHDR of the credal set $\mathsf{CRED}(y^{n},\psi)$. ∎
@@
Hence, $\psi_{\theta}\in\mathscr{F}$. Notice that consonance, i.e. $\sup_{y\in Y}\pi(y,y^{n})=1$, does not automatically hold in this example. To enforce it, it is enough to consider the normalized conformal transducer $\tilde{\pi}(y,y^{n})\coloneqq\frac{\pi(y,y^{n})}{\sup_{y\in Y}\pi(y,y^{n})}$ (Cella and Martin,, 2021, Section 7), which is still upper semicontinuous because the denominator is a strictly positive constant for fixed $y^{n}$.
@@
    Heteroscedastic conformal regression: If $\varphi_{\theta}$ is the penultimate layer of a regression network, (9) yields locally adaptive, yet still valid, prediction intervals.
Hence, this choice of $\psi_{\theta}$ simultaneously fulfills joint continuity and practical usefulness in modern ML workflows. $\triangle$

### v3

@@ Abstract.
Conformal prediction (CP) is an Uncertainty Representation technique that delivers finite-sample calibrated prediction regions for any underlying Machine Learning model. Its status as an Uncertainty Quantification (UQ) tool, though, has remained conceptually opaque: While Conformal Prediction Regions (CPRs) give an ordinal representation of uncertainty (larger regions typically indicate higher uncertainty), they lack the capability to cardinally quantify it (twice as large regions do not imply twice the uncertainty). We adopt a category-theoretic approach to CP—framing it as a morphism, embedded in a commuting diagram, of two newly-defined categories—that brings us three joys. First, we show that—under minimal assumptions—CP is intrinsically a UQ mechanism, that is, its cardinal UQ capabilities are a structural feature of the method. Second, we demonstrate that CP bridges the Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning. Finally, we show that a CPR is the image of a covariant functor. This observation is relevant to AI privacy: It implies that privacy noise added locally does not break the global coverage guarantee.
@@ 1. Introduction
For example, Stehlík, (2012) and Song, (2024) survey how to frame statistical learning models into a CT framework. Shiebler et al., (2021) do the same for gradient-based and probabilistic methods, and invariant and equivariant learning. Furthermore, Cornish, (2025) expands on the latter by studying the problem of symmetrizing a neural network along a group homomorphism. That is, given two groups $H$ and $G$, they study a procedure that converts $H$-equivariant neural networks to $G$-equivariant ones. More in general, CT approaches to probability and statistics (which play a pivotal role in modern-day ML) are active fields of study, see e.g. Fritz et al., (2025); Bohinen and Perrone, (2025); Ensarguet and Perrone, (2024); Ackerman et al., (2024); Matache et al., (2022), and are widely discussed in the nLab and in the n-Category Café.
@@
    • As a byproduct, the user will now be able to see how much the choice of non-conformity measure $\psi$ affects (different types of) uncertainty in the analysis.
@@
  • Privacy by Design. Seeing the Conformal Prediction Region as the image of a covariant functor directly benefits AI privacy (Fang et al.,, 2025): (i) Locally added privacy noise leaves coverage guarantees intact, and (ii) In federated learning (Liu et al.,, 2024), each agent need only share a summary object (the credal set $\mathcal{M}$, introduced in the following paragraphs) instead of raw data.
@@
To make the paper largely self-contained, we give the necessary background on CP in Section 2. There, we also provide some concepts from the Imprecise Probability (IP) literature. IPs (Walley,, 1991; Augustin et al.,, 2014; Troffaes and de Cooman,, 2014) are mathematical models that allow the scholar to reason in the presence of uncertainty (as for classical probabilities), and ambiguity (that is, ignorance, full or partial, about the “true” data generating process). [^4] For more modern references, see e.g. (Caprio and Gong,, 2023; Caprio and Mukherjee,, 2023; Caprio and Seidenfeld,, 2023; Caprio et al., 2024b, ; Caprio,, 2024, 2025). Their applications to ML form a lively subject field, see e.g. Caprio et al., 2024a ; Sale et al., (2023, 2024); Caprio et al., 2024c ; Lu et al., (2024); Dutta et al., (2024); Chau et al., (2025). Functional Analysis and Category Theory notions are brushed up in Appendices C and D, respectively. The structure of the paper can be summarized as follows.
@@
In Section 3, we introduce category $\mathbf{UHCont}$. Its objects are topological spaces, and its morphisms are upper hemicontinuous correspondences. Under minimal assumptions on the state space $Y$, on the choice of the credibility level $\alpha$, and on the class $\mathscr{F}$ of non-conformity scores $\psi$, we show that CP is a morphism of $\mathbf{UHCont}$. Viewing CP as a morphism rather than a stand-alone algorithm lets ML engineers plug it into functional-style libraries (e.g. JAX or PyTorch Functorch) and obtain automatic batching, JIT compilation, and differentiability of the score map, while the coverage guarantee remains mathematically intact.
CP is also a morphism of category $\mathbf{WMeas}_{\text{uc}}$ (which we briefly introduce in Section 3, and explore in depth in Appendix A to keep the paper easily readable) whose objects are compact Polish spaces and whose morphisms are weakly measurable, uniformly compact-valued correspondences. We study both $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$ because continuity and measurability are fundamental—but generally non-equivalent—concepts in functional analysis.
@@
Not only is the CP correspondence a morphism of these categories, but it is also part of a commuting diagram in both. As we shall see (in particular in Example 1 and in the discussion thereafter), this discovery implies the intrinsic (cardinal) UQ capabilities of CP.
@@
In Section 4, we show that, under minimal assumption, a diagram that codifies the way Bayesian CP (Fong and Holmes,, 2021), classical CP, and CP-informed imprecise probability theory generate an $\alpha$-level prediction region, commutes asymptotically. This is a strong indicator that CP forms a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning (Martin,, 2022).
@@
Section 5 expands on category-theoretic Imprecise Probability theory (Liell-Cock and Staton,, 2024). We formalize the notion of Imprecise Highest Density Region (Coolen,, 1992)—that can be thought of as a “robust version” of a confidence interval which also accounts for distribution uncertainty—as a (covariant) functor $\mathsf{IHDR}_{\alpha}$, $\alpha\in[0,1]$, between the category of closed and convex sets of finitely additive probabilities on $Y$, and that of subsets of $Y$, both with inclusion morphisms. In turn, the Conformal Prediction Region (CPR) can be seen as the image $\mathsf{IHDR}_{\alpha}(\mathcal{M})$ of such a functor at the set of probabilities $\mathcal{M}$ associated with the CP procedure (Cella and Martin,, 2022). This has e.g. the following promising implication. If we have many sources producing credal sets $\mathcal{M}_{j}$, then privacy-preserving operation applied locally to each credal set $\mathcal{M}_{j}$ propagates automatically through the functor. The resulting prediction regions, then, still satisfy the global coverage guarantee while only the (possibly privatized) summary objects $\mathcal{M}_{j}$, never the raw data, need be shared.
@@ 2.1. Full Conformal Prediction
This section is an expanded reproduction of Caprio et al., 2025a (, Section 2.1). It is based on Cella and Martin, (2022, Section 4.1), who summarize work by Shafer and Vovk, (2008) and Vovk et al., (2005) on Transductive (or Full) Conformal Prediction.
@@
for all $i\in\{1,\ldots,n+1\}$, where $y^{n+1}_{-i}=y^{n+1}\setminus\{y_{i}\}$ and $\psi:Y^{n}\times Y\rightarrow\mathbb{R}$ is a fixed function that is invariant to permutations in its first vector argument. Function $\psi$, which is called a non-conformity measure, is constructed in such as way that $\psi_{i}(y^{n+1})$ is small if and only if $y_{i}$ agrees with—i.e. is “close to”—a prediction based on the data $y^{n+1}_{-i}$. Or, the other way around, large values $\psi_{i}(y^{n+1})$ suggest that the observation $y_{i}$ is “strange” and does not conform to the rest of the data $y^{n+1}_{-i}$. The key idea is to define $\psi_{i}(y^{n+1})$ in a way that allows us to compare $y_{i}$ to a suitable summary of $y^{n+1}_{-i}$, e.g. $\psi_{i}(y^{n+1})=|\text{mean}(y^{n+1}_{-i})-y_{i}|$, for all $i\in\{1,\ldots,n+1\}$. Notice that transformation $\mathcal{Y}^{n+1}\rightarrow T^{n+1}$ preserves exchangeability.
@@ 2.2. Imprecise Probabilities
Let $(Y,\Sigma_{Y})$ be a measurable space, and call $\Delta_{Y}$ the space of finitely additive probabilities on $Y$, endowed with the weak topology. A credal set $\mathcal{M}$ is a weak-closed and convex element of $\Delta_{Y}$ (Levi,, 1980). Throughout the paper, we will denote by $\mathscr{C}\subset 2^{\Delta_{Y}}$ the space of nonempty weak-closed and convex subsets of $\Delta_{Y}$. For any nonempty credal set $\mathcal{M}\in\mathscr{C}$, define the upper probability as the upper envelope of $\mathcal{M}$,
@@
Even in the presence of such ambiguity, we are able to define (a version of) a confidence interval. Fix a credibility level $\alpha\in[0,1]$. Then, the $\alpha$-level Imprecise Highest Density Region (IHDR) of credal set $\mathcal{M}$ is the set $B_{\mathcal{M}}^{\alpha}\subset Y$ such that $\underline{P}(B_{\mathcal{M}}^{\alpha})=1-\alpha$, and its “size is minimal” (Coolen,, 1992). More formally (Caprio et al., 2025a, , Equation (11)),
@@
We also consider a slight modification of $\mathsf{CRED}$, that will significantly simplify our proving endeavors in Sections 3 and 4, and in Appendix A. We denote it by $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\text{, }\forall A\in\Sigma_{Y}\text{ closed}\}$. As we can see, $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\supseteq\mathsf{CRED}(y^{n},\psi)=\mathcal{M}(\overline{\Pi})$, because it imposes less constraints on the $P$’s: They only have to be setwise dominated on closed sets. The upper probability $\overline{\Pi}$ and its conjugate lower probability $\underline{\Pi}$ (the only characteristics of the credal sets that we use) stay the same for both $\mathsf{CRED}$ and $\mathsf{CRED}_{\text{cl}}$: This justifies our simplifying choice. To avoid notation complications, in the remainder of this work we do not distinguish between $\mathsf{CRED}(y^{n},\psi)$ and $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)$.
@@ 3. Conformal Prediction as an Upper Hemicontinuous Morphism
  • Morphisms: Upper hemicontinuous correspondences $\Phi:X\rightrightarrows Y$ (see Definition 25 in Appendix C).
@@
We verify the axioms of a category. Recall that the composition $\Psi\circ\Phi$ of two correspondences $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ is given by $(\Psi\circ\Phi)(x)\coloneqq\bigcup_{y\in\Phi(x)}\Psi(y)$.
@@
(3) Closure under composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be upper hemicontinuous correspondences. By Aliprantis and Border, (2006, Theorem 17.23), the composition $\Psi\circ\Phi$ is also upper hemicontinuous. Thus, the composition of morphisms is again a morphism.
@@
    $\Delta_{Y}$ is endowed with the weak topology, and $\mathscr{C}\subset 2^{\Delta_{Y}}$ is endowed with the Vietoris topology (see Definition 28 in Appendix C).
@@
Another important lemma that sets the table for our main result is the following. To improve readability, we defer its proof to Appendix B.
@@
Let us also emphasize that we only use $\mathsf{IHDR}_{\alpha}$ through its restriction to the image set $\operatorname{Im}(\mathsf{CRED})\subset\mathscr{C}$ (which is trivially an object of $\mathbf{UHCont}$). We denote it by $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}\equiv\mathsf{IHDR}_{\alpha}\mid_{\operatorname{Im}(\mathsf{CRED})}:\operatorname{Im}(\mathsf{CRED})\rightrightarrows Y$. We do not claim that the stand alone correspondence $\mathsf{IHDR}_{\alpha}:\mathscr{C}\rightrightarrows Y$ is upper hemicontinuous on all of $\mathscr{C}$; This is not needed for the diagram to commute inside $\mathbf{UHCont}$, and may fail without additional assumptions.
@@
Fix any $n\in\mathbb{N}$. We first show that $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ have a closed graph.
(i) $\kappa_{\alpha}$. By (ii) and its definition, the conformal transducer $\pi(\cdot,y^{n},\psi):Y\to[0,1]$ is upper semicontinuous. Pick any $\alpha$ that avoids the grid $S_{n+1}$, i.e. $\alpha\in\Lambda\coloneqq[0,1]\setminus S_{n+1}$. Let
@@
This is a superlevel set of $\pi(\cdot,y^{n},\psi)$, and since $\pi$ is upper semicontinuous, $G$ is closed. In turn, $\operatorname{Gr}(\kappa_{\alpha})=G$ is closed.
@@
Weak-continuity of $P\mapsto P(A)$ and upper semicontinuity of the right-hand side (see Lemma 4) make $G_{A}$ closed. Because $\operatorname{Gr}(\mathsf{CRED})=\bigcap_{A\in\Sigma_{Y}\text{, }A\text{ closed}}G_{A}$, the graph is closed.
(iii) $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$. Under the consonance assumption, for each $(y^{n},\psi)$, the credal set $\mathsf{CRED}(y^{n},\psi)$ is generated by the possibility $\overline{\Pi}_{(y^{n},\psi)}(A)=\sup_{y\in A}\pi_{\psi}(y,y^{n})$, $A\in\Sigma_{Y}$. Recall that, for any $\alpha\notin S_{n+1}$, we defined $\beta(\alpha)\coloneqq\min\{\tfrac{k}{n+1}\in S_{n+1}:\tfrac{k}{n+1}>\alpha\}$. Then, by (Caprio et al., 2025a, , Proposition 5), we have that
$\mathsf{IHDR}_{\alpha}\big{(}\mathsf{CRED}(y^{n},\psi)\big{)}=\Bigl{\{}y\in Y:\pi_{\psi}(y,y^{n})\geq\beta(\alpha)\Bigr{\}}.$
By Lemma 4, $y\mapsto\pi_{\psi}(y,y^{n})$ is upper semicontinuous on $Y$, so the $\alpha$-cut $\{\pi_{\psi}(\cdot,y^{n})\geq\beta(\alpha)\}$ is closed in $Y$. Moreover, since $(y,y^{n},\psi)\mapsto\pi_{\psi}(y,y^{n})$ is upper semicontinuous, the graph
$\operatorname{Gr}\bigl{(}\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}\bigr{)}=\Bigl{\{}\bigl{(}y^{n},\psi,y\bigr{)}\in Y^{n}\times\mathscr{F}\times Y:\pi_{\psi}(y,y^{n})\geq\beta(\alpha)\Bigr{\}}$
is closed. Therefore, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ has closed graph.
Notice also that $Y$ is compact Hausdorff by (i), and that $\Delta_{Y}$ is compact by (iv) and Banach-Alaoglu, [^6] To be precise, let $B_{b}(Y,\Sigma_{Y})$ be the Banach space of bounded $\Sigma_{Y}$-measurable functions with the sup norm. Its (isometric) dual is $ba(\Sigma_{Y})$, the space of bounded finitely additive signed measures (charges) on $\Sigma_{Y}$, with the variation norm. Equip $ba(\Sigma_{Y})$ with the weak topology. Then, $\Delta_{Y}$ is a weak-closed subset of the weak-compact unit ball of $ba(\Sigma_{Y})$, hence $\Delta_{Y}$ is compact Hausdorff in this topology. and Hausdorff because it is a subset of a dual Banach space equipped with its weak topology.
It follows that $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ are upper hemicontinuous by Lemma 3. In turn, they are morphisms of $\mathbf{UHCont}$.
$Y$ is a topological space from (i); $\operatorname{Im}(\mathsf{CRED})$ and $\mathscr{C}$ are topological spaces by (iv); $Y^{n}\times\mathscr{F}$ is a topological space in the product topology by (ii). Hence, these spaces are all objects of $\mathbf{UHCont}$.
The fact that the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}$, then, is an immediate consequence of Caprio et al., 2025a (, Proposition 5). It states that, for all $\alpha\in[0,1]$, $y^{n}\in Y^{n}$, and $\psi\in\mathscr{F}$, the Conformal Prediction Region $\kappa_{\alpha}(y^{n},\psi)$ corresponds to the $\alpha$-level IHDR of the credal set $\mathsf{CRED}(y^{n},\psi)$, in formulas $\mathsf{IHDR}_{\alpha}(\mathsf{CRED}(y^{n},\psi))$. Of course, here we restrict to $\alpha\in[0,1]\setminus S_{n+1}$. ∎
@@
Hence, $\psi_{\theta}\in\mathscr{F}$. Notice that consonance, i.e. $\sup_{y\in Y}\pi(y,y^{n})=1$, does not automatically hold in this example. To enforce it, it is enough to consider the normalized conformal transducer $\tilde{\pi}(y,y^{n})\coloneqq\frac{\pi(y,y^{n})}{\sup_{y\in Y}\pi(y,y^{n})}$ (Cella and Martin,, 2021, Section 7), which is still upper semicontinuous because the denominator is a strictly positive constant for fixed $y^{n}$, and it returns the same CP Region (after the corresponding rescaling of $\alpha$).
@@
    Heteroscedastic conformal regression: If $\varphi_{\theta}$ is the penultimate layer of a regression network, (9) yields locally adaptive, yet still valid, prediction sets.
Hence, this choice of $\psi_{\theta}$ simultaneously fulfills joint continuity and practical usefulness in modern ML workflows. $\diamond$